OpenStudy (austinl):

Consider a tank used in certain hydrodynamic experiments. After one experiment the tank contains 200L of dye solution with a concentration of 1 g/L. To prepare for the next experiment, the tank is to be rinsed with fresh water flowing in at a rate of 2 L/min, the well-stirred solution flowing out at the same rate. Find the time that will elapse before the concentration of dye in the tank reaches 1% of its original value.

OpenStudy (austinl):

@Bella♥ Take a crack at it if you like :)

OpenStudy (austinl):

Ok, I am gonna try and talk myself through it on here.

OpenStudy (austinl):

\(\large{\dfrac{dQ}{dt}=rate~in-rate~out}\)

OpenStudy (anonymous):

Lets see hmm....

OpenStudy (austinl):

The concentration of the liquid entering the tank is \(\dfrac{0g}{liter}\) at a rate of \(\dfrac{2liters}{min}\)..... so it goes in at \(\dfrac{0g}{min}\).

OpenStudy (austinl):

I am properly confused by this problem.

OpenStudy (austinl):

@terenzreignz If you have any comments, feel very very free to share :P

OpenStudy (anonymous):

Mee too! LAME!!! okay okay lets see if I can work thi out on paper hold upp!

terenzreignz (terenzreignz):

I was hoping you wouldn't say that... hang on, let's see if I can somehow forget that it's a Friday xD

OpenStudy (anonymous):

200L * 1g/L = 200g of dye at the beginning. 2L of the solution are flowing out and 2L of fresh water is flowing in, so the tank is losing (1g/L)(-2L/min) = -2g/min of dye. 1% of 200g = 2g, so to reach that dilution the tank has lost 198g of dye. This will take (1min/2g)*(198g) = 99 min = 1 hour 39 minutes. I think is how u do this! Hope I am right ahah I am not sure!

terenzreignz (terenzreignz):

this fails to take into account the 'well-stirred' -ness of the solution... It only works if the drains drain a solution with 1g/L concentration every time, which isn't true...

OpenStudy (austinl):

Unfortunately no, that isn't how you are supposed to do it I believe. The answer in the back of my book is \(\approx460.5~min\)

OpenStudy (anonymous):

Lame! Sorry than :/

OpenStudy (austinl):

Thank you for trying :)

OpenStudy (anonymous):

¯t = −100 ln(1/100) m = 100 ln 100 m. If you use a calculator, you see that is is about 460.51701859880916 m, or a bit less than 8 hours.

OpenStudy (anonymous):

The volume of liquid in the tank is always 200 L. Let Q(t) be the amount of die in the tank, measured in grams. The concentration of die in the tank at any given time is given by Q(t)/200. We have Q(0) = 200 L × 1 g/L = 200 g. The rate of change of Q(t) is given by dQ(t) dt = rate of die going into tank − rate of dye leaving tank = 0 g/L × 2 L/m − Q(t) 200 g/L × 2 L/min = − 1 100 Q(t) g/m Solving for Q(t) gives Q(t) = 200e−t/100 g. There will be 2 g of die in the tank when it is 1% of the original value. So we want to find the time ¯t when 2 g = 200 e−¯t/100 g, or The answer up above :)

OpenStudy (anonymous):

I think you will find this answer is right :P

OpenStudy (austinl):

Don't mean to look the proverbial "gift horse" in the mouth or anything, but you should really learn to use \(\LaTeX\). It makes your responses much, MUCH easier to decipher. As is, I am having some problems following your work in the above response.

OpenStudy (austinl):

@terenzreignz I am gonna link in help that I received on another site. If you could help explain a later part in the answer of ChrisLT521 it would be appreciated. http://mathhelpboards.com/differential-equations-17/modeling-first-order-equation-6296.html#post28617

terenzreignz (terenzreignz):

last part... starts from where?

OpenStudy (austinl):

\(\text{"Once you solve this simple ODE, you're then left with finding t such that x(t)=.01x(0)"}\) This is confusing me and I feel it shouldn't be.

OpenStudy (austinl):

Once you solve this simple ODE, you're then left with finding t such that x(t)=.01x(0

terenzreignz (terenzreignz):

HAVE you solved the ODE?

OpenStudy (austinl):

\(\dfrac{dx}{dt}=-\dfrac{x}{100}\)? I am sorry, that is what I meant, how would I go about that? I feel really, really dumb right now.

terenzreignz (terenzreignz):

Probably because it's Friday...rearrange... \[\Large \frac{dx}x = -\frac{dt}{100}\] Integrate both sides.

OpenStudy (austinl):

\(\int\dfrac{1}{x}dx = \int -\dfrac{1}{100}dt\) Right?

terenzreignz (terenzreignz):

Yup.

OpenStudy (austinl):

\(\ln|x|=-\dfrac{t}{100}+C\)?

terenzreignz (terenzreignz):

I prefer it like this: \[\Large \ln(x) = -\frac{t}{100}+ \ln(C)\] ln(C) is just as arbitrary a constant, and it makes this easier: Raise e to both sides of the equation, you get \[\Huge x = Ce^{-\frac{t}{100}}\]

terenzreignz (terenzreignz):

Now, you need to find C.

terenzreignz (terenzreignz):

ideas? :)

OpenStudy (austinl):

Not really unfortunately. I am not sure where the "Initial Value" comes in.... God I feel like a dunce right now.

terenzreignz (terenzreignz):

Well, you do know what x should be when t = 0. Before mixing in any fresh water, the concentration was 1g/L and there were 200L

OpenStudy (austinl):

200g yes?

terenzreignz (terenzreignz):

Yes.... so x(0) = 200. Work it :D

OpenStudy (austinl):

derp...

terenzreignz (terenzreignz):

So C = 200. \[\Huge x = 200 e^{-\frac{t}{100}}\]

terenzreignz (terenzreignz):

Now, for the concentration to be 1% of its original concentration (200), x has to be equal to 1% of 200, which is 2. \[\Huge \color{blue}2 = 200e^{-\frac{t}{100}}\]

OpenStudy (austinl):

I got it!!!! YEEEEEESSS!

terenzreignz (terenzreignz):

Awesome ^_^

OpenStudy (austinl):

I hope every problem isn't this hard on this assignment.